OpenStudy (anonymous):

A cone and a cylinder both of circular bases have the same height and the same volume. Find the ratio of there radii.

OpenStudy (anonymous):

The ratio of radii from cone to cylinder is √3:1

OpenStudy (anonymous):

can you show me how you get it?

OpenStudy (anonymous):

Sure! let R be the radius of cone and r be the radius of the cylinder so 1/3pie R*R*h=pie*r*r*h this gives R*R= 3*r*r r/R = 1/root 3